OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2

terenzreignz (terenzreignz):

Rearranging the first equation gives us : y = 8 - 4x Rearranging the second equation give us : 2y = x - 2 Taking the rearranged first equation, and multiplying 2 to both sides, we get : 2y = 16 - 8x Since 2y = 16 - 8x and 2y = x - 2, then 16 - 8x = x - 2 -9x = - 18, therefore x = 2